camlewis00:

Jason's savings account has a balance of $2179. After 5 years, what will the amount of interest be at 6% compounded quarterly?

bm717:

The formula for those equations will be A=P(1+r)^t.

bm717:

And you would have to insert the numbers in there. A=2179(1+(0.06/4)^5*10

bm717:

P is the original amount of money you had, thats why that balance is there. You divide the 4 by the interest because it is QUARTERLY. And sorry, I made a mistake, but the time is supposed to be 5*4.

bm717:

A=2179(1+(0.06/4)^5*4

bm717:

If i did this right, the answer would be 2934.80

sillybilly123:

yep